OpenStudy (ashleyy_smith):

Translate this sentence into an equation. The difference of Raj's age and 13 is 42 Use the variable r to represent Raj's age.

OpenStudy (oryzae):

r - 13 = 42

OpenStudy (ashleyy_smith):

can you help me with two more @Oryzae ? (: ~~~ 12 more than Jose's age is 75. ~~~

OpenStudy (oryzae):

Using J to represent Jose's age; J + 12 = 75. "more than" indicates addition, and 75 would be the total sum.

OpenStudy (ashleyy_smith):

What about if it says, "68 decreased by Donnie's savings is 10." ? @Oryzae

OpenStudy (oryzae):

That's a bit weirdly worded... but I'm assuming it means "decreased by" means subtracting D, Donnie's savings. 68 - D = 10.

OpenStudy (ashleyy_smith):

Thanks @Oryzae (: <3